OpenStudy (anonymous):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form. A. y = 2x B. y = -2x C. y = 1/2 x D. y = -1/2 x

OpenStudy (mysesshou):

Let k be the slope: \[k = \frac{change In Y}{changeIn X}= \frac{Y_2-Y_1}{X_2-X_1}= \frac{-2-0}{1-0}= \frac{-2}{1}=-2\]